OpenStudy (anonymous):

What is the volume of a rectangular based pyramid with a height of fifteen meters and base dimensions of 4 meters and 3 meters? sixty cubic meters one hundred eighty cubic meters two hundred thirty-four cubic meters five hundred forty cubic meters

OpenStudy (anonymous):

Do you know the equation f or the volume of a rectangular pyramid?

OpenStudy (anonymous):



OpenStudy (anonymous):

So we are given l = 3, w = 4, and h = 15 Do you know what to do with that information now?

OpenStudy (anonymous):

Here is a picture to show.

OpenStudy (anonymous):

Do I multiply?

OpenStudy (anonymous):

Why yes, so a rectangle based pyramid has a volume of 1/3 times the base times the width times the height. This can be shown by the equation: V = 1/3*l*w*h Or V = (l*w*h)/3

OpenStudy (anonymous):

Since you have length, width, and height (4, 3, 15) respectively, you can calculate it!

OpenStudy (anonymous):

Remember all of these are in units of meters. And that by multiplying meters three times will give your answer in meters^3. V = (l * w * h)/3 V = (4 * 3 * 15)/3 V = 60 V = 60 m^3